OpenStudy (anonymous):

How many sides does a regular polygon have if each exterior angle measures 72o?

OpenStudy (anonymous):

exterior angle = 360/number of sides. So 72 = 360/n , n = 360/72 = 5 , so 5 sides which is a regular pentagon

OpenStudy (anonymous):

Sum of the exterior angles of a polygon is 360, so if each ext. angle is 72, divide 360 by 72 to find number of sides. 360/72=5

OpenStudy (anonymous):

oh. Thank you guys so much. Very helpful.